A 2.0-kg brick has the dimensions 7.5 cm x 15 cm x 30 cm. Find the pressures 

exerted by the brick on a table when it is resting on its various faces


1
Expert's answer
2021-06-21T11:42:03-0400

(a) Let's first find the pressure the brick exerts when standing on its largest face:


"P=\dfrac{F}{A}=\dfrac{mg}{ab}=\dfrac{2\ kg\cdot9.8\ \dfrac{m}{s^2}}{0.15\ m\cdot0.3\ m}=435.5\ Pa."

(b) Let's find the pressure the brick exerts when standing on two other faces:


"P=\dfrac{F}{A}=\dfrac{mg}{ac}=\dfrac{2\ kg\cdot9.8\ \dfrac{m}{s^2}}{0.15\ m\cdot0.075\ m}=1742\ Pa,""P=\dfrac{F}{A}=\dfrac{mg}{bc}=\dfrac{2\ kg\cdot9.8\ \dfrac{m}{s^2}}{0.3\ m\cdot0.075\ m}=871\ Pa."